Codes, Lattices, Modular Forms and Vertex Operator Algebras

Octorber 2 – 4, 2000 Faculty of Science, Yamagata University

Organizers: Michio Ozeki (Yamagata University)

Eiichi Bannai (Kyushu University) Masaaki Harada (Yamagata University)

PROGRAM

Octorber 2 Monday 10:00–10:10 Opening 10:10–11:00 Rudolf Scharlau (Universität Dortmund, Germany)

Classifying lattices using modular forms 11:10–12:00 Robin Chapman (University of Exeter, UK)

Double circulant constructions of lattices 13:30–14:20 Ming-Chang Kang (National Taiwan University, Taiwan)

Lüroth’s Theorem revisited 14:30–15:10 Steven T. Dougherty (University of Scranton, USA)

Shadows of codes over rings of order 4 15:30–16:20 Michel Deza (Ecole Normale Supérieure, France)

Rank 1 forms, closed zones and laminae 16:30–17:00 Hiroki Aoki (Ritsumeikan University)

Estimation of the dimension of modular forms by using Jacobi forms

Octorber 3 Tuesday

9:30–10:20 Tomoyoshi Ibukiyama (Osaka University)

A general theory of differential operators on Siegel modular forms and special polynomials 10:30–11:20 Boris Venkov (Steklov Mathematical Institute in St. Petersburg, Russia)

On Siegel modular forms of weight 12 11:30–12:00 Nobuhiko Ishida (Osaka Institute of Technology)

and Noburo Ishii (Osaka Prefecture University) Generators and defining equation of the modular function field of the group Γ_{1}(N)

13:30–14:20 Masahiko Miyamoto (Tsukuba University)

VOAs and modular forms of rational number weights 14:30–15:00 Haesuk Kim (Pohang University of Science and Technology, Korea)

Codes over ℤ_{2m} and Jacobi forms of genus n 15:20–16:10 Hidenori Katsurada (Muroran Institute of Technology)

Dirichlet series for Siegel modular forms and quadratic forms 16:20–17:00 Yoshiyuki Kitaoka (Meijo University)

Distribution of units of a cubic field with negative discriminant

Octorber 4 Wednesday

9:30–10:20 Siegfried Böcherer (Universität Mannheim, Germany)

Modular forms of high vanishing order and applications to quadratic forms 10:30–11:20 Patrick Solé (CNRS, ESSI, France)

Rankin-Cohen brackets and transvectants 11:30–12:00 Jon-Lark Kim (University of Illinois at Chicago, USA)

On additive self-dual codes over GF(4) and lattices 13:20–14:10 Hiromichi Yamada (Hitotsubashi University)

Lattice VOAs and Virasoro algebras 14:20–14:50 Masaaki Kitazume (Chiba University)

Classification of 12-dimensional even unimodular Gaussian lattices 15:00–15:20 Manabu Oura (Kyushu University)

A theta relation in genus 4 15:20–15:40 Keisuke Shiromoto (Kumamoto University)

Maximum distance codes over finite rings 15:40–16:00 Masaaki Harada (Yamagata University)

ℤ_{22}-code construction of the Leech lattice